A contract has the following data: Target Cost = 300K Target Fee = 15K Max Fee = 20K Min Fee = 5K Share Ratio = 80/20 Actual Cost = 360K What is the final price that would be paid to the contractor?
Answer #1

Actual cost = 360 K

Target fee = 15 K

Contractor's share of overrun = (360-300) * 20% = 12K

Contractor's fee = Target fee - Contractor's share of overrun = 15k -12 k = 3k , however minimum fee is 5k.

Final price paid to contractor = Actual cost + Contractor's fee = 360k + 5k = 365K